# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v1

@@ 3.3 Computation of marginal encoder density
$\mathrm{d}z_{t}$ $=$ $\left(f(z_{t},t)-\frac{1}{2}g(t)^{2}s(z_{t},x,t|\theta)\right)\mathrm{d}t$ (8)
$\triangleq$ $f^{PF}(z_{t},x,t|\theta)\mathrm{d}t$
This alignment holds true if $s(z_{t},x,t|\theta)$ effectively approximates the score function $\nabla_{z_{t}}\log p_{E}(z_{t}|x)$ after training.
As a consequence, given $x$ and $z_{0}$, we can numerically solve the ODE to trace the path $z^{PF}_{t}$, initiated with $z^{PF}_{0}=z_{0}$. The computation of the marginal encoder density $p_{E}(z_{0}|x)$ adheres to the principles of neural ODE (Chen et al., 2018), expressed as:
$\log p_{E}(z_{0}|x)=\log p_{E}(z_{T}^{PF})+\int_{0}^{T}\mathrm{tr}\left(\frac{\partial{f}^{PF}(z^{PF}_{t},x,t|\theta)}{\partial z^{PF}_{t}}\right)\mathrm{~{}d}t.$ (9)
Based on the aforementioned findings, we can efficiently estimate the normalizing constant $Z$ and utilize importance sampling techniques to further correct modeling errors in EDG. (See Appendix B.)
3.4 Decoding with generalized Hamiltonian dynamics
In our investigation, we observed that EDG performs poorly for multimodal distributions, primarily due to the mode-seeking behavior of the KL divergence. To address this limitation, we propose a decoder based on generalized Hamiltonian dynamics (GHD) (Song et al., 2017; Levy et al., 2017), summarized in Algorithm 1, which significantly enhances the sampling performance for multimodal distributions.
Within the algorithm, we initially stochastically generate a sample $y$ using the random variable $\zeta$ (refer to Line 1 of the algorithm). Subsequently, we iteratively update the sample using random velocities $v_{1},\ldots,v_{K}$ and Generalized Hamiltonian Dynamics (see Line 5). Finally, we utilize discretized Brownian dynamics with a trainable step size to design the decoder density of $x$ (refer to Line 9). Notably, the parameters of $A$, $T_{v}$, $Q_{v}$, $\epsilon$, and $\eta$ are all trainable, and $\epsilon_{0}$ serves as a hyperparameter. Our numerical experiments suggest that setting $\epsilon_{0}$ to a small positive value can improve the stability of the algorithm.
Algorithm 1 GHD based decoder
2:Let $y:=a(\zeta;\phi)$.
3:for $k=1,\ldots,K$ do
4:     for $j=1,\ldots,J$ do
5:         Let $s:=\frac{(k-1)J+j-1}{KJ}$.
6:         Updata $(y,v_{i})$ by
$v_{k}$ ::=:= $v_{k}-\frac{\epsilon(s;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),s;\phi)}+T_{v}(y,\nabla U(y),s;\phi)\right)$
$y$ ::=:= $y+\epsilon(s;\phi)\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},s;\phi)}+T_{y}(v_{k},s;\phi)\right)$
$v_{k}$ ::=:= $v_{k}-\frac{\epsilon(s;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),s;\phi)}+T_{v}(y,\nabla U(y),s;\phi)\right)$
7:     end for
8:     Let $v_{k}:=-v_{k}$.
9:end for
10:Let $\mu=y-e^{\epsilon_{0}\eta(y;\phi)}\nabla U(y)$ and $\Sigma=2e^{\epsilon_{0}\eta(y;\phi)}I$.
11:return $x\sim\mathcal{N}(\mu,\Sigma)$.
In the experiments of this paper, unless explicitly stated otherwise, EDG will utilize a GHD based decoder.
@@ 4 Experiments
We empirically evaluate EDG across a diverse range of energy functions. Initially, we present the results obtained from a set of two-dimensional distributions. Subsequently, we demonstrate the performance of EDG in the context of Bayesian logistic regression. All the experimental details are provided in Appendix D.
@@
w/o GHD
HMC
EDG
w/ GHD
Figure 2: Density plots for 2D energy function. For the generation of reference samples, please refer to Appendix D. We generate $500,000$ samples for each method and plot the histogram.
2D energy function
Firstly, we compare our model with HMC (Neal et al., 2011), VAE and BG (Noé et al., 2019) on several synthetic 2D energy functions: MoG2(i) (Mixture of two isotropic Gaussians with the same $\sigma^{2}=0.5$ or different variance $\sigma_{1}^{2}=1.5,\sigma_{2}^{2}=0.3$ and centroids separated by distance $10$), MoG6 (Mixture of six isotropic Gaussians with variance $\sigma^{2}=0.1$), MoG9 (Mixture of nine isotropic Gaussians with variance $\sigma^{2}=0.3$) Ring, Ring5. For HMC, we conducted 500 independent runs, each providing 1,000 samples after a burn-in phase of 1,000 steps. The VAE architecture closely mirrors that of EDG, utilizing the loss defined in (1). Here, the decoder incorporates a prior distribution for $z_{0}$ denoted by $p_{D}(z_{0})\sim\mathcal{N}(0,I)$, and both $p_{D}(x|z_{0})$ and $p_{E}(z_{0}|x)$ are modeled as Gaussian distributions. Additionally, multilayer perceptrons (MLPs) are employed to compute the means and variances of $p_{D}(x|z_{0})$ and $p_{E}(z_{0}|x)$. The BG employs RealNVP, comprising 6 affine coupling layers, where the scaling and shifting operations for each layer are executed by MLPs with two hidden layers, and each layer of MLPs consists of 256 hidden units. We draw the histogram of samples for visual inspection in Fig. 2. In addition, for EDG, we also compared the sampling performance between a GHD based decoder (EDG w/ GHD) and a conventional decoder (EDG w/o GHD). The sampling results are summarized in Fig. 2 and Table 1. It can be seen that the EDG with GHD based decoder outperforms other methods.
The Maximum Mean Discrepancy (MMD) of 2D energy tasks.
Mog2 Mog2(i) Mog6 Mog9 Ring Ring5
VAE $1.86$ $1.62$ $2.57$ $0.17$ $0.11$ $0.36$
BG $1.90$ $1.63$ $2.64$ $0.07$ $0.05$ $0.18$
HMC $\mathbf{0.01}$ $1.56$ $0.02$ $0.04$ $\mathbf{0.01}$ $0.82$
EDG w/o GHD $0.06$ $1.30$ $0.04$ $0.05$ $0.09$ $0.05$
EDG w/ GHD $\mathbf{0.01}$ $\mathbf{1.14}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $\mathbf{0.03}$
Table 1: The maximum mean discrepancy between the samples generated by each generator and the reference samples. Details on the calculation of discrepancy can be found in Appendix D.
@@
EDG $\mathbf{83.15\pm 1.01}$ $\mathbf{91.63\pm 0.47}$ $\mathbf{78.94\pm 0.92}$ $\mathbf{78.11\pm 0.70}$ $\mathbf{87.56\pm 2.23}$ $\mathbf{94.35\pm 0.76}$
Table 2: Classification accuracy and AUC results for Bayesian logistic regression tasks. The experiments utilize consistent training and test data partitions, where the HMC step size is set to $0.01$, and EDG utilizes the GHD based decoder. Average accuracy and AUC values, accompanied by their respective standard deviations, are computed across $32$ independent experiments for all datasets.
HMC VAE BG EDG
Acc $49.88\pm 3.32$ $50.99\pm 2.67$ $50.75\pm 3.78$ $\mathbf{70.13\pm 2.13}$
Table 3: Classification accuracy on the test dataset of Coverstype. The reported values represent averages and standard deviations of accuracy over 32 independent experiments. The GHD based decoder is used in EDG.
Bayesian Logistic Regression
In the subsequent experiments, we highlight the efficacy of EDG in the context of Bayesian logistic regression, particularly when dealing with a posterior distribution residing in a higher-dimensional space. In this scenario, we tackle a binary classification problem with labels $L=\{0,1\}$ and high-dimensional features $D$. The classifier’s output is defined as
$p(L=1|D,x)=\mathrm{softmax}(w^{\top}D+b),$
where $x=(w,b)$. Drawing samples $x^{1},\ldots,x^{N}$ from the posterior distribution
based on the training set $\mathcal{D}_{\mathrm{train}}$, where the prior distribution $p(x)$ is a standard Gaussian distribution. For $L,D$ in the test dataset, $p(L|D,\mathcal{D}_{\mathrm{train}})$ is approximated as $\sum_{n}p(L|D,x^{n})$. We conduct experiments on three datasets: Australian (AU, 15 covariates), German (GE, 25 covariates), and Heart (HE, 14 covariates) (Dua & Graff, 2017), evaluating accuracy rate (ACC) and Area Under the Curve (AUC) on the test subset. Remarkably, as illustrated in Table 2, EDG consistently achieves the highest accuracy and AUC performance.
We extend our analysis to the binary Covertype dataset comprising 581,012 data points and 54 features. The posterior of the classifier parameters follows a hierarchical Bayesian model (matching the model in Section 5 of (Liu & Wang, 2016)), with $x$ denoting the combination of classifier parameters and the hyperparameter in the hierarchical Bayesian model. To enhance computational efficiency, in VAE, BG, and EDG, $\log\pi(x)$ is unbiasedly approximated during training as
@@
where $\mathcal{B}$ is a random mini-batch. For HMC, the exact posterior density is calculated. As indicated by the results in Table 3, EDG consistently outperforms alternative methods.
@@ 5 Conclusion
Looking forward, our primary focus will be on extending the application of EDG to construct generative models for large-scale physical and chemical systems, such as proteins (Zheng et al., 2023; Jing et al., 2023).
Appendix A Proof of (5)
@@
Appendix B Normalizing constant approximation and importance sampling via marginal encoder density
Eq. (1) provides a lower bound for $\log Z$ as follows:
$\log Z\geq-\mathbb{E}_{p_{D}(x,z_{0}|\phi)}\left[\log\frac{p_{D}(z_{0})p_{D}(x|z_{0},\phi)}{p_{E}(z_{0}|x,\theta)}+U(x)\right].$
The tightness of this bound is achieved when $p_{D}(x,z_{0}|\phi)$ closely approximates $p_{E}(x,z_{0}|\theta)=\pi(x)\cdot p_{E}(z_{0}|x,\theta)$. Based on (9), we can employ an ODE solver to efficiently compute $p_{E}(z_{0}|x,\theta)$, thereby enabling the calculation of a variational approximation for $\log Z$.
Furthermore, leveraging the value of $p_{E}(z_{0}|x,\theta)$ obtained from the ODE solver, we can unbiasedly estimate the value of $\mathbb{E}{\pi}[O(x)]$ through importance sampling. Let $x^{1},\ldots,x^{N}$ represent $N$ samples generated by the decoder in EDG with latent variables $z{0}^{1},\ldots,z_{0}^{N}\sim\mathcal{N}(0,I)$. The estimation is given by:
$\mathbb{E}_{\pi(x)}[O(x)]$ $=$ $\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0})}\left[O(x)\frac{\pi(x)p_{E}(z_{0}|x)}{p_{D}(z_{0})p_{D}(x|z_{0})}\right],$
$\approx$ $\frac{1}{N}\sum_{n}O(x^{n})\frac{\pi(x^{n})p_{E}(z_{0}^{n}|x^{n})}{p_{D}(z_{0}^{n})p_{D}(x^{n}|z_{0}^{n})}.$
Appendix C Importance weighting for accurate time integral

### v2

@@ 3.2 Loss function
$+\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\Bigg{[}\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\nabla_{z_{t}}\cdot s\left(z_{t},x,t;\theta\right)$
$\hskip 170.71652pt+\left\|\nabla_{z_{t}}\log p_{D}(z_{t})\right\|^{2}\Bigg{]}\mathrm{d}t.$
By utilizing the importance-sampled integration and Hutchinson estimator, we can obtain an equivalent expression of $\mathcal{L}(\theta,\phi)$, which can be efficiently and unbiasedly estimated by Monte Carlo random sampling from $p_{D}$ without numerical solutions to SDEs:
where
$p(\epsilon)$ is the Rademacher distribution with $\mathbb{E}[\epsilon]=0,\mathrm{cov}(\epsilon)=I$, $p(t)$ is a proposal distribution of $t\in[0,T]$, and the weighting function $\lambda(t)=\frac{g(t)^{2}}{2p(t)}$. All neural networks involved in EDG can then be trained by minimizing the loss function (7) using stochastic gradient descent.
In our experiments, to mitigate the variance inherent in the Monte Carlo approximation of the loss function, we adopted the strategy from [49] to dynamically adjust the time proposal distribution $p_{D}(t)$, based on a historical buffer that stores recent loss values related to $t$ in the r.h.s of (7). For more details, please refer to D.
3.3 Sample reweighting
After training, we can use the decoder $p_{D}(x|z_{0})$ to generate samples and compute various statistics of the target distribution $\pi(x)$. For example, for a quantity of interest $O:\mathbb{R}^{d}\to\mathbb{R}$, we can draw $N$ augmented samples $\{(x^{n},z_{0}^{n})\}_{n=1}^{N}$ from $p_{D}(z_{0})p_{D}(x|z_{0})$ and estimate the expectation $\mathbb{E}_{\pi(x)}[O(x)]$ as follows:
$\mathbb{E}_{\pi(x)}[O(x)]\approx\frac{1}{N}\sum_{n}O(x^{n}).$
However, due to model errors, this estimation can be systematically biased. To address this, we can apply importance sampling, using $p_{D}(x,z_{0})=p_{D}(z_{0})p_{D}(x|z_{0})$ as the proposal distribution and $p_{E}(x,z_{0})=\pi(x)p_{E}(z_{0}|x)$ as the augmented target distribution. We can then assign each sample $(x,z_{0})$ generated by the decoder an unnormalized weight:
$w(x,z_{0})=\frac{\exp(-U(x))p_{E}(z_{0}|x)}{p_{D}(z_{0})p_{D}(x|z_{0})}\propto\frac{\pi(x)p_{E}(z_{0}|x)}{p_{D}(z_{0})p_{D}(x|z_{0})}$ (8)
and obtain a consistent estimate of $\mathbb{E}_{\pi(x)}[O(x)]$ as:
$\mathbb{E}_{\pi(x)}[O(x)]\approx\frac{\sum_{n}w(x^{n},z_{0}^{n})O(x^{n})}{\sum_{n}w(x^{n},z_{0}^{n})},$
where the estimation error approaches zero as $N\to\infty$.
In addition, the weight function $w$ can also be used to estimate the normalizing constant $Z$, which is a crucial task in many applications, such as Bayesian model selection in statistics and free energy estimation in statistical physics. Based on (1), we have:
$\log Z\geq-\mathbb{E}_{p_{D}(x,z_{0})}\left[\log w(x,z_{0})\right].$
Here, the lower bound can also be estimated using samples from the decoder, and the tightness of this bound is achieved when $p_{D}(x,z_{0})=p_{E}(x,z_{0})$.
The main difficulty in above computations lies in the intractability of the marginal decoder density $p_{E}(z_{0}|x)$ when calculating the weight function. To overcome this, following the conclusion from Sec. 4.3 in [37], we construct the following probability flow ordinary differential equation (ODE):
$\mathrm{d}z_{t}=\left(f(z_{t},t)-\frac{1}{2}g(t)^{2}s(z_{t},x,t;\theta)\right)\mathrm{d}t,$
with the boundary condition $z_{T}\sim p_{E}(z_{T})$. If $s(z_{t},x,t;\theta)$ accurately approximates the score function $\nabla_{z_{t}}\log p_{E}(z_{t}|x)$ after training, the conditional distribution of $z_{t}|x$ given by the ODE will match that of the encoding process for each $t\in[0,1]$. Consequently, we can employ the neural ODE method [50] to efficiently compute $p_{E}(z_{0}|x)$.
@@ 4 Experiments
We conduct an empirical evaluation of EDG across a diverse range of energy functions. First, we present the results obtained from a set of two-dimensional distributions. Next, we showcase the performance of EDG in the context of Bayesian Logistic regression. Last, we apply EDG to an Ising model. All the experimental details are provided in E. Additionally, we perform an ablation study to verify the effectiveness of each module in EDG. Please refer to F for more information.
In order to demonstrate the superiority of our model, we compare EDG with the following sampling methods:
  • 1.
    Vanilla Hamiltonian Monte Carlo method [8], denoted as V-HMC.
  • 2.
    L2HMC [13], a GHD-based MCMC method with a trainable proposal distribution model.
  • 3.
    Boltzmann Generator (BG) [18], a VI method where the surrogate distribution is modeled using RealNVP [51].
  • 4.
    Neural Renormalization Group (NeuralRG) [17], a method similar to BG and designed specifically for the Ising model. In this section, NeuralRG is only used for experiments of the Ising model.
  • 5.
    Path Integral Sampler (PIS) [39], a diffusion-based sampling model via numerical simulation of an SDE.
Table 1: The Maximum Mean Discrepancy (MMD) between the samples generated by each generator and the reference samples. Details on the calculation of discrepancy can be found in E.
Mog2 Mog2(i) Mog6 Mog9 Ring Ring5
V-HMC $\mathbf{0.01}$ $1.56$ $0.02$ $0.04$ $\mathbf{0.01}$ $\mathbf{0.01}$
L2HMC $0.04$ $0.94$ $\mathbf{0.01}$ $0.03$ $0.02$ $\mathbf{0.01}$
BG $1.90$ $1.63$ $2.64$ $0.07$ $0.05$ $0.18$
PIS $\mathbf{0.01}$ $1.66$ $\mathbf{0.01}$ $0.42$ $\mathbf{0.01}$ $0.78$
EDG $\mathbf{0.01}$ $\mathbf{0.50}$ $\mathbf{0.01}$ $\mathbf{0.02}$ $\mathbf{0.01}$ $0.02$
@@
Figure 2: Density plots for 2D energy function. For the generation of reference samples, please refer to E. We generate $500,000$ samples for each method and plot the histogram.
2D energy function Firstly, we compare our model with the other models on several synthetic 2D energy functions: MoG2(i) (Mixture of two isotropic Gaussians with the same $\sigma^{2}=0.5$ or different variance $\sigma_{1}^{2}=1.5,\sigma_{2}^{2}=0.3$ and centroids separated by distance $10$), MoG6 (Mixture of six isotropic Gaussians with variance $\sigma^{2}=0.1$), MoG9 (Mixture of nine isotropic Gaussians with variance $\sigma^{2}=0.3$), Ring, Ring5 (energy functions can be seen in [12]). We present the histogram of samples for visual inspection in Fig. 2, and Table 1 summarizes the sampling errors. As shown, EDG delivers higher-quality samples compared to the other methods. To elucidate the function of each component in EDG, we compare our model with vanilla VAE in F.
Bayesian Logistic Regression In the subsequent experiments, we highlight the efficacy of EDG in the context of Bayesian logistic regression, particularly when dealing with a posterior distribution residing in a higher-dimensional space. In this scenario, we tackle a binary classification problem with labels $L=\{0,1\}$ and high-dimensional features $D$. The classifier’s output is defined as
$p(L=1|D,x)=\mathrm{softmax}(w^{\top}D+b),$
where $x=(w,b)$. We aim to draw samples $x^{1},\ldots,x^{N}$ from the posterior distribution
based on the training set $\mathcal{D}_{\mathrm{train}}$, where the prior distribution $p(x)$ is a standard Gaussian distribution. Then, for a given $D$, the conditional distribution $p(L|D,\mathcal{D}_{\mathrm{train}})$ can be approximated as $\sum_{n}p(L|D,x^{n})$. We conduct experiments on three datasets: Australian (AU, 15 covariates), German (GE, 25 covariates), and Heart (HE, 14 covariates) [52], evaluating accuracy rate (ACC) and Area Under the Curve (AUC) on the test subset. Remarkably, as illustrated in Table 2, EDG consistently achieves the highest accuracy and AUC performance.
Table 2: Classification accuracy and AUC results for Bayesian logistic regression tasks. The experiments utilize consistent training and test data partitions, where the HMC step size is set to $0.01$. Average accuracy and AUC values, accompanied by their respective standard deviations, are computed across $32$ independent experiments for all datasets.
@@
PIS $81.64\pm 2.63$ $91.23\pm 0.67$ $71.90\pm 3.17$ $71.67\pm 4.52$ $83.24\pm 3.95$ $91.68\pm 2.78$
EDG $\mathbf{84.96\pm 1.67}$ $\mathbf{92.82\pm 0.69}$ $\mathbf{79.40\pm 1.74}$ $\mathbf{82.79\pm 1.46}$ $\mathbf{88.02\pm 3.90}$ $\mathbf{95.10\pm 1.23}$
We extend our analysis to the binary Covertype dataset comprising 581,012 data points and 54 features. The posterior of the classifier parameters follows a hierarchical Bayesian model (see Sec. 5 of [27]), with $x$ denoting the combination of classifier parameters and the hyperparameter in the hierarchical Bayesian model. To enhance computational efficiency, in BG and EDG, $\log\pi(x)$ is unbiasedly approximated during training as
@@
where $\mathcal{B}$ is a random mini-batch. For V-HMC and L2HMC, the exact posterior density is calculated. As indicated by the results in Table 3, EDG consistently outperforms alternative methods.
Table 3: Classification accuracy on the test dataset of Coverstype. The reported values represent averages and standard deviations of accuracy over 32 independent experiments.
V-HMC L2HMC BG PIS EDG
Acc $49.88\pm 3.32$ $51.51\pm 3.46$ $50.75\pm 3.78$ $50.59\pm 2.94$ $\mathbf{70.13\pm 2.13}$
Table 4: The estimation of $\log Z_{\mathrm{Ising}}$ in 2D ising model with a dimension of 256 ($16\times 16$) is obtained by the method described in Sec. 3.3. We utilize a batch size of $n=256$ to estimate the mean and applies the central limit theorem to calculate the standard deviation of the mean of the statistic as $\mathrm{std}/\sqrt{n}$.
$\log Z_{\mathrm{Ising}}$ NeuralRG PIS EDG
$T=2.0$ $260\pm 0.13$ $210\pm 0.43$ $\mathbf{270\pm 0.18}$
$T=2.1$ $250\pm 0.14$ $208\pm 0.41$ $\mathbf{255\pm 0.19}$
$T=2.2$ $239\pm 0.16$ $210\pm 0.39$ $\mathbf{252\pm 0.17}$
$T=2.3$ $231\pm 0.15$ $214\pm 0.37$ $\mathbf{233\pm 0.17}$
$T=2.4$ $\mathbf{225\pm 0.17}$ $212\pm 0.37$ $\mathbf{225\pm 0.15}$
$T=2.5$ $219\pm 0.17$ $202\pm 0.37$ $\mathbf{221\pm 0.14}$
$T=2.6$ $\mathbf{216\pm 0.18}$ $181\pm 0.40$ $214\pm 0.14$
$T=2.7$ $\mathbf{212\pm 0.18}$ $189\pm 0.36$ $\mathbf{212\pm 0.14}$
Ising model Finally, we verify the performance of EDG on the 2-dimensional Ising model [17], a mathematical model of ferromagnetism in statistical mechanics. To ensure the continuity of physical variables, we employ a continuous relaxations trick [53] to transform discrete variables into continuous auxiliary variables with the target distribution:
$\pi(\boldsymbol{x})=\exp\left(-\frac{1}{2}\boldsymbol{x}^{T}\left(K(T)+\alpha I\right)^{-1}\boldsymbol{x}\right)\times\prod_{i=1}^{N}\cosh\left(x_{i}\right),$
where $K$ is an $N\times N$ symmetric matrix depending on the temperature $T$, $\alpha$ is a constant guaranteeing $K+\alpha I$ to be positive. For the corresponding discrete Ising variables $\boldsymbol{s}=\{1,-1\}^{\otimes N}$, one can directly obtain discrete samples according to $\pi(\boldsymbol{s}|\boldsymbol{x})=\prod_{i}\left(1+e^{-2s_{i}x_{i}}\right)^{-1}$. When there is no external magnetic field and each spin can only interact with its neighboring spins, $K$ is defined as $\sum_{<ij>}s_{i}s_{j}/T$, with the nears neighboring sum $<ij>$. Consequently, the normalizing constant of the continuous relaxations system is given by $\log Z=\log Z_{\text{Ising }}+\frac{1}{2}\ln\operatorname{det}(K+\alpha I)-\frac{N}{2}[\ln(2/\pi)-\alpha]$ [17]. Additionally, using the method described in Sec. 3.3, we provide lower bound estimates of $\log Z_{\text{Ising}}$ at different temperatures for samples generated by NeuralRG, PIS and EDG. Since these are lower bound estimates, larger values indicate more accurate results. As seen in Table 4, EDG provides the most accurate estimates of $\log Z$ across most temperature ranges. The generated states at different temperatures are displayed in Fig. 3.
Figure 3: The generated states at different temperatures from $T=2.0$ to $T=2.7$ by EDG with a dimensionality of 256 ($16\times 16$), where the latent variable $z_{0}$ remains fixed. As the temperature rises, the model’s state tends progressively toward disorder.
@@ 5 Conclusion
In consideration of its powerful generation ability and unrestrained network design theoretically, there is still room for further exploration. We can design specific network architectures for different tasks and find the most elaborate one. Nevertheless, it is the first attemp to design a one-shot generator assisted by diffusion models.
In future, our primary focus will be on extending the application of EDG to construct generative models for large-scale physical and chemical systems, such as proteins [54, 55].
Acknowledgements
The first and third authors are supported by the NSF of China (under grant number 12171367). The second author is supported by the NSF of China (under grant numbers 92270115, 12071301) and the Shanghai Municipal Science and Technology Commission (No. 20JC1412500), and Henan Academy of Sciences. The last author was supported by the NSF of China (No. 12288201), the Strategic Priority Research Program of Chinese Academy of Sciences (Grant No. XDA25010404), the National Key R&D Program of China (2020YFA0712000), the Youth Innovation Promotion Association, CAS, and Henan Academy of Sciences.
Appendix A Proof of the independence between $z_{T}$ and $x$
In the decoding process, if $z_{T}$ is independent of $z_{0}$, we have
$p_{D}(z_{T}|x)$ $=$ $\int p_{D}(z_{T}|z_{0})p_{D}(z_{0}|x)\mathrm{d}z_{0}$
$=$ $p_{D}(z_{T})\cdot\int p_{D}(z_{0}|x)\mathrm{d}z_{0}$
$=$ $p_{D}(z_{T}).$
Appendix B Proof of (5)
@@
Considering integration by parts, $D_{KL}(p_{D}||p_{E})$ in Eq.(6) has the following equivalent expression
$\mathbb{E}_{p_{D}}\left[s\left(z_{t},x\right)^{\top}\nabla_{z_{t}}\log p_{D}(z_{t}|z_{0})|x,z_{0}\right]$ $=$ $\int p_{D}(z_{t}|z_{0})s\left(z_{t},x\right)^{\top}\nabla_{z_{t}}\log p_{D}(z_{t}|z_{0})\mathrm{d}z_{t}$
$=$ $\int s\left(z_{t},x\right)^{\top}\nabla_{z_{t}}p_{D}(z_{t}|z_{0})\mathrm{d}z_{t}$
$=$ $\int\mathrm{div}_{z_{t}}\left(p_{D}(z_{t}|z_{0})s\left(z_{t},x\right)\right)\mathrm{d}z_{t}$
$-\int p_{D}(z_{t}|z_{0})\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)\mathrm{d}z_{t}$
$=$ $-\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)|z_{0}\right]$
$\Rightarrow\mathbb{E}_{p_{D}}\left[s\left(z_{t},x\right)^{\top}\nabla_{z_{t}}\log p_{D}(z_{t}|z_{0})\right]$ $=$ $-\mathbb{E}_{p_{D}}\left[\mathrm{tr}\left(\frac{\partial s(z_{t},x)}{\partial z_{t}}\right)\right]$
Substituting the above equation to Eq.(6), we have
$D_{KL}\left(p_{D}(x,z_{[\cdot]})||p_{E}(x,z_{[\cdot]})\right)$ $=$ $\mathbb{E}_{p_{D}}\left[\log\frac{p_{D}(x|z_{0};\phi)}{\pi(x)}\right]$
$+\mathbb{E}_{p_{D}}\left[\frac{g^{2}(t)}{2}\left(\left\|\nabla\log p_{D}(z_{t})\right\|^{2}+\left\|s\left(z_{t},x,t|\theta\right)\right\|^{2}\right.\right.$
$\quad\quad\quad\quad\ \quad\quad\left.\left.-2s\left(z_{t},x,t|\theta\right)^{\top}\nabla\log p_{D}(z_{t}|z_{0})\right)\right].$
Appendix C Implementation of decoder
The detailed implementation of the GHD based decoder is summarized in Algorithm 1. Within the algorithm, we initially stochastically generate a sample $y$ using the random variable $\zeta$ (refer to Line 1 of the algorithm). Subsequently, we iteratively update the sample using random velocities $v_{1},\ldots,v_{K}$ and GHD (see Line 5). Finally, we utilize discretized Brownian dynamics with a trainable step size to design the decoder density of $x$ (refer to Line 9). Here, $\epsilon(l;\phi)$, $Q_{v}$, $Q_{y}$, $T_{v}$ and $T_{y}$ are all neural networks composed of three-layer MLPs. $\epsilon_{0}$ serves as a hyperparameter, and our numerical experiments suggest that setting $\epsilon_{0}$ to a small positive value can improve the stability of the algorithm.
Algorithm 1 GHD based decoder
2:Let $y:=a(\zeta;\phi)$.
3:for $k=1,\ldots,K$ do
4:     for $j=1,\ldots,J$ do
5:         Let $l:=\mathrm{leap\ frog\ step\ of\ GHD}$.
6:         Updata $(y,v_{i})$ by
$v_{k}$ ::=:= $v_{k}-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right)$
$y$ ::=:= $y+\epsilon(l;\phi)\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},l;\phi)}+T_{y}(v_{k},l;\phi)\right)$
$v_{k}$ ::=:= $v_{k}-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right)$
7:     end for
8:     Let $v_{k}:=-v_{k}$.
9:end for
10:Let $\mu=y-e^{\epsilon_{0}\eta(y;\phi^{\prime})}\nabla U(y)$ and $\Sigma=2e^{\epsilon_{0}\eta(y;\phi^{\prime})}I$.
11:return $x\sim\mathcal{N}(\mu,\Sigma)$.
Appendix D Importance weighting for accurate time integral